A ball thrown up vertically returns to the thrower after 6 s. Find
       the velocity with which it was thrown up,

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

u = ?
v = 0
g = –9.8 m/s^2(thrown upward)
Total time = 6 s (to go up and down)
∴ Time for upward journey = 6/2 = 3s
(a) v = u + gt
0 = u = (–9.8) × 3
u = 29.4 m/s

Thanks